### v1

# Ink and Individuality: Crafting a Personalised Narrative in the Age of LLMs

Abstract.
Individuality and personalization comprise the distinctive characteristics that make each writer unique and influence their words in order to effectively engage readers while conveying authenticity. However, our growing reliance on LLM-based writing assistants risks compromising our creativity and individuality over time. We often overlook the negative impacts of this trend on our creativity and uniqueness, despite the possible consequences. This study investigates these concerns by performing a brief survey to explore different perspectives and concepts, as well as trying to understand people’s viewpoints, in conjunction with past studies in the area. Addressing these issues is essential for improving human-computer interaction systems and enhancing writing assistants for personalization and individuality.
copyright: none
1. Introduction
The exceptional performance of large language models (LLMs) across a wide range of applications makes them increasingly well-liked in academia and industry (Chang et al., 2023). Especially with chatbots such as Microsoft Copilot, Gemini, and ChatGPT, which have the transformative capability of producing polished and well-crafted stories, essays and assignments akin to those written by a professional novelist or a writer (Chung et al., 2022; Yuan et al., 2022). Not to mention, other writing tools such as Quillbot and Grammarly have made writing a straightforward, time-efficient task now. With recent works such as (Kim et al., 2023) and (Yeh et al., 2024), we are at a junction where personalization is of high significance with LLM usage for writing. This raises a question: Are users concerned about the level of personalization in their writings with the emergence of AI-assisted writing tools? In this paper, we explore (i) the importance of personalization and individuality of writing and (ii) personalization in the age of LLMs. In this study of these issues, we try to understand the perspective of individuals through a thoroughly conducted survey, along with previous research work done in this domain. We aim to contribute to the development of personalization in writing with LLMs along with its significance and try to highlight whether individuals are affected by it. It is important to highlight the personal touch of an individual when writing content, for example, crafting personal statement essays for university applications or cover letters for jobs, which are important factors in whether someone gets accepted or rejected.
2. Personalization and Individuality in Writing
Personalization in writing means customizing the content to meet the needs and preferences of the audience, while still maintaining the author’s distinct voice and perspective. This includes adjusting the language, tone, and examples to connect with various readers. On the other hand, individuality refers to the unique traits and style that distinguish each writer, influencing the tone and content of their work to convey genuineness and effectively engage readers. Striking a balance between personalization and individuality enables writers to create captivating and relatable content that resonates with diverse audiences, while also showcasing their own unique identity.
From writing alone, most of the time, it is possible to understand the inner feelings, characteristics and emotions of the writer. Therefore, it is safe to say every writer has a unique method of writing, which establishes their writing style. While creativity is a big factor in enhancing the quality of writing, some programs tend to develop rigid rules, restricting the level of individuality and personalization. Therefore, critical thinking abilities are further improved if one is encouraged to add a personal touch to one’s writing (Krueger, 2001). Most notably, for educational purposes, personalization plays a big role in ensuring teachers are giving proper constructive feedback catered to each student (Wang and Li, 2022) and students are putting their own thoughts forward rather than blatantly plagiarizing (Milano et al., 2023). The recent emergence of AI-assisted tools for writing task completion has observed millions of users employing them for writing purposes. Spellcheckers (Wang et al., 2021) were among the first computational writing aids. Further advances include techniques for determining writing style (Sterman et al., 2020), ideas for brainstorming support tools (Gero and Chilton, 2019a, b), and theories of cognitive writing for creating writing frameworks (Hui et al., 2018). With all these tools, it may be observed that the writing style of individuals might become similar or lose their individual rhythm, making it monotonous and giving content traits that are suggestive of AI-driven creation and synonymous with automated generation.
Figure 1. Survey Statistics: Frequency of LLM Assistance in Different Types of Writing
Figure 2. Survey Statistics: View on Personalization and LLMs
3. Personalization in the Age of LLMs
LLMs show potential for helping users with a range of creative writing activities, including screenplays (Mirowski et al., 2023) and short stories (Yang et al., 2022; Yuan et al., 2022). LLMs have become useful tools for helping users without a background in computing by responding directly to user commands after being aligned to output content more in accordance with human preferences (Ouyang et al., 2022). Nevertheless, modern LLMs continue to lack a robust fact-checking method, which could result in errors, mainly when writing non-fiction (Jiang et al., 2020). From what appears to be harmless information, they can create detrimental details (Gehman et al., 2020; Perez et al., [n. d.]), putting minorities in danger of being misrepresented. Sociopolitical and environmental challenges also come up as ethical issues (Bender et al., 2021). Although prompt engineering is becoming more popular to handle these issues (Bach et al., 2022; Liu et al., 2023), non-technical users may find it intimidating due to its intricacy. Another concern of LLM usage for writing is prominent in academia is plagiarism (Park, 2017), where education bodies must decide how to react: ban it or incorporate it in the teaching and examining systems (Milano et al., 2023). Padmakumar and He (Padmakumar and He, 2023) conduct an experiment to contrast writing between model-free assisted writing, feedback-tuned writing (InstructGPT), and utilizing a basic model (GPT-3). InstructGPT’s results indicate that it considerably reduces diversity, increasing the similarity of texts written by different writers. This decrease results from the model producing content that is less varied, even when user contributions stay the same.
4. Problems and Perspectives
To investigate the role of LLMs in content personalization and individuality, we conducted a brief survey with 35 participants. The demographic profile of the participants are available in Section A. The responses that we received show different perspectives and concerns in this area. Figure 1 shows the frequency of LLM assistance in three types of writing: personal writing (letters, messages and birthday wishes), formal writing (CVs and applications), and academic writing (reports, assignments and homework).
The term ”Personal Touch” in writing, in this paper, refers to the unique voice, style, and perspective that an individual writer brings to their work. It encompasses a distinct tone, relatable personal anecdotes or experiences, clear opinions and viewpoints, preferred writing style and word choice, as well as the ability to forge an emotional connection with the reader. The personal touch is what allows a writer’s personality and humanity to shine through their words, distinguishing their writing from others and making it feel more authentic, intimate, and engaging.
The survey reveals widespread use of LLMs across various writing contexts. Personal writing sees the least utilization (14.3% people never use LLMs) due to a desire to maintain original emotional expression. Conversely, LLM usage is nearly ubiquitous in formal writing, with only around 3% abstaining, and even more prevalent in academic writing, where it’s becoming a standard practice. While content personalization remains crucial in all contexts, its significance appears to vary, with academic settings showing a trend towards increased reliance on LLMs (28.6% use them everytime), hinting at a potential evolution in educational practices.
So, do LLMs negatively affect personalization and individuality? Our survey data reveals that a majority (71.4%) value personal touches in writing (Figure 2 (a)), and many believe LLMs can capture their personal touch (Figure 2 (b)). However, despite this, 60% of respondents feel their personal touch is lost when using LLMs (Figure 2 (c)), suggesting a discrepancy between perception and reality regarding the impact of LLMs on personalization and individuality.
So, the answer is yes. Extensive use of LLMs in writing diminishes personalization and individuality by homogenizing content. These AI systems often prioritize efficiency and conformity over unique expression, leading to a loss of the author’s distinct voice and style. Furthermore, reliance on LLMs can discourage individuals from cultivating their writing skills and developing their own creative approaches, resulting in a dependency that hampers personal growth and self-expression. Additionally, the standardized nature of LLM-generated content may dilute the emotional resonance and authenticity that characterize truly personalized writing, detracting from its impact on readers.
Although participants acknowledge the significance of personal touches in writing and think LLMs can capture these touches, a significant number of the respondents still feel that their personal touch is lost when using LLMs. This discrepancy raises questions about possible limitations in current LLMs’ ability to accurately replicate individual style and expression. One of the reasons for believing LLMs can capture personal touches is similarity in tone or structure. However, the participants actually feel their personal touch is lost because nuances in expression or unique writing quirks are not faithfully replicated, leading to a sense of impersonal or generic output. Another possible explanation is that participants perceive a decrease in their writing or thinking ability as they put less effort, leading them to feel concerned that their personal touch is being lost.
Furthermore, these gaps emphasize the need for more research and development to address the challenges associated with maintaining personalization and individuality in the age of AI writing assistants. It also raises questions about the effectiveness of LLMs in preserving the authenticity and uniqueness of personal writing, suggesting that reliance on these tools may lead to a perceived loss of personalization and individuality despite attempts to integrate personal touches.
5. Why does losing our individuality matter?
From our survey, we have discovered that given three modes of writing: personal, academic and formal, they prefer to leverage LLMs to generate content. One interesting finding is that for personal writing, we have the same number of individuals using LLMs sometimes and most of the time. But what we can conclude from our findings is that the majority of users aren’t bothered too much about LLMs not being the greatest agents of personalized writing. As long as informative information is captured, they are satisfied.
But, lack of personalization with extensive LLM usage can lead to generic, impersonal content that fails to engage readers on a deeper level or resonate with their specific interests and needs (Bekker, 2024). Similarly, a lack of individuality can result in homogenized writing that lacks uniqueness and fails to showcase the author’s voice and perspective. Over time, this may lead to decreased reader interest, reduced credibility, and a loss of connection between the writer and audience. Therefore, raising awareness about the importance of personalization and individuality in writing is crucial for fostering more meaningful and impactful communication.
6. Discussions on Limitations and Future Works
We acknowledge that there are some shortcomings in the survey questionnaire regarding quantifying the definitions of options and word choices for certain questions. As this is an ongoing endeavor, we aim to rectify these issues in future surveys and data collections. We also acknowledge that exploring the correlation between perception towards the need and loss of personal touch with each of these domains would be insightful. We will certainly consider incorporating this aspect into future analyses. Further, we notice a gap regarding the lack of clarity on what each participant perceives as a ’personal touch.’ Indeed, understanding individual perspectives on this matter could significantly enhance the depth and context of our analysis, particularly in relation to questions concerning the capability of LLMs to capture personal touch and the perceived loss of personal touch when using LLMs. We will clarify these issues to the further survey participants.
7. conclusion
This study investigates the different obstacles and perspectives of writing with personalization and individuality in the age of LLMs. According to our survey, we are currently heavily reliant on LLMs for writing any kind of content, from personal to format to academic; and this heavy use of LLMs is harming our individuality, but we are not too concerned about it. This study connects our survey with literature to understand individual perspectives and contribute to the development of personalised writing with LLMs.
Below is a concise summary of both the data-supported findings regarding participants’ motives and some speculative insights:
  • We investigate whether writers are concerned about adding personalized content to their writings and if LLMs can capture it. We find out most writers have claimed that personal touch is important in writing, but LLMs fail to capture their persona and nature for content writing.
  • We also discover that writers are not too burdened by LLMs’ lack of ability to capture their individuality and personal touch as long as the main gist is identified, which is very concerning.
  • We emphasize the need for features or modes in current LLMs that will help obtain the personal touch in writing to embrace an individual’s uniqueness. Future discussions and research on this area can focus on this aspect.
Appendix A Survey Participant Demographics
The demographic data of survey participants presents a demographic consisting primarily of university students aged 18-24 (94.3%), with a smaller proportion (5.7%) in the 24-30 age group. The gender distribution is skewed towards males (65.7%), with females making up 34.3% of the group. Notably, the overwhelming majority (88.6%) come from a Science background, while a small percentage (8.6%) are from Arts backgrounds. Significantly, most individuals in this demographic rated their technology awareness as high, scoring between 3 and 5 on a scale where 5 represents a ”tech nerd” and 1 indicates being unaware of technology. This suggests a cohort of tech-savvy individuals, likely pursuing STEM fields in their university studies.